OpenStudy (anonymous):

In an AP the 1st term is 2 and the sum of the 1st 5 terms is one-fourth of the next 5 terms.Show that 20th term is -112.

OpenStudy (ash2326):

Let the AP be a1, a2, a3....an First term is a1=2 Sum of the first five terms=\(a1+a2+a3+a4+a5=5\frac{(a1+a5)}{2}\) do you get this @JOYAL?

OpenStudy (anonymous):

@ash2326 yes, can u complete it?

OpenStudy (amistre64):

hard to parse: "the sum of the 1st 5 terms is one-fourth of the next 5 terms" but that could just be me

OpenStudy (ash2326):

Sum of the next five terms= \[a6+a7+a8+a9+a10+a11=5\frac{(a6+a11)}{2}\] @joyal now we have the relation \[\frac{1}{4}\times5\frac{(a1+a5)}{2}= \frac{(a6+a11)}{2}\] could you plugin the values of a5, a6, and a11 in terms of a1 and the comon difference d

OpenStudy (amistre64):

you might have your 1/4 on the wrong side if a is 1/4 of b; then, a = b/4

OpenStudy (ash2326):

oops @amistre64 Sorry ! \[5\frac{(a1+a5)}{2}\ \\ \ \ \ = \frac{1}{4}\times\frac{(a6+a11)}{2}\]

OpenStudy (anonymous):

@amistre64 so u r saying that the question is wrong?

OpenStudy (amistre64):

im saying the question is hard to read. but i follow ashes concept, and do not disagree that it is a plausible rendering :)

OpenStudy (anonymous):

@amistre64 @ash2326 thanks for ur help..

OpenStudy (amistre64):

youre welcome, so far :)

OpenStudy (amistre64):

\[4*5*\frac{(2+a_5)}{2}= 5*\frac{(a_6+a_{11})}{2}\] if we take that and reduce it we get \[2(2+a_5)=a_6+a_{11}\]

OpenStudy (amistre64):

which i cant see working for us at the moment ....

OpenStudy (amistre64):

if we rework the AP in terms of another one say \[b_1=(2+a_2+a_3+a_4+a_5)\] \[b_2=(2+a_2+a_3+a_4+a_5+d)/4\] this might be promising

OpenStudy (amistre64):

after some working out, that thought went nowhere if we focus on the partials a1 = 2+0d a2 = 2+1d a3 = 2+2d a4 = 2+3d a5 = 2+4d ---------- 10+10d a6 = 2+5d a7 = 2+6d a8 = 2+7d a9 = 2+8d a10 = 2+9d ----------- 10+35d 10+10d = (10+35d)/4 40+40d = 10+35d 5d = 50 d = 10 does; 2+19(10) = -112 ? since it doesnt, it means that we are most likely reading the problem incorrectly

OpenStudy (amistre64):

pfft, forgot how to subtract :) 10-40 = -30; -30/5 = -6 2+19(-6) = -112 ?